# Antifragility of RIS-assisted Communication Systems under Jamming Attacks

### v1

@@ IV-E Modulation and Code Adaptation
Since $(n,k)$ Reed-Solomon code can correct up to $t=(n-k)/2$ symbol errors, choosing the optimal code depends on the post-demodulation error profile. Thus, the measured SNR is used to compute the bit error rate (BER) for each adaptive modulation scheme, which is then converted into the residual symbol error rate $P^{\mathrm{RS}}_{\mathrm{res}}$ following the formulation in [15].
@@
IV-F Antifragile gain
We denote the throughput in the baseline scenario by $T_{L}$ and under jamming attack by $T_{J}$ respectively. The optimal Reed-Solomon code is selected according to Eq. (18), which ensures that all symbol errors within each block are corrected. The throughput is then defined based on the code rate $R_{C}$, bandwidth $B$, modulation order $M$ and SNR index $i$ (Eq.  17):
@@ IV-F Antifragile gain
The system demonstrates antifragile gain only if its throughput during jamming surpasses its baseline, jammer-free throughput. To evaluate this gain across different SNRs, the jamming-to-signal ratio (JSR) is defined. It represents the ratio of jamming power $P_{J}$ to legitimate signal power $P_{L}$:
@@ V Numerical Evaluation
V-A Assumptions
[…]
V-B Source-Aware Eavesdropping
Figure 2 compares the three attacks with a fixed code rate. Without spatial orthogonality the jammer and desired signal overlap for half the frame; the transmitter shortens its burst (Sec. IV-B, Eq. 14), reducing payload and throughput (Eq. 19). Thus, antifragile gains appear only after raising the modulation order, i.e., at JSR = 3 dB for DRFM and 15 dB for AS. With full orthogonality (Eq. 12) and a higher baseline SNR, throughput exceeds the jammer, free reference from about -5 dB, DRFM allowing for the largest boost due to its high mutual information with the desired waveform. Figure 3 introduces adaptive RS coding based on the Eqs. 17-18 and a higher receiver SNR. Variable code rates enable both an earlier appearance and a larger magnitude of antifragile gain by pairing higher-order modulations (increasing $M$ in Eq. 18) with moderately reduced code rates ($R_{C}$ in Eq. 18), thereby enhancing error-correction capability and elevating the peak throughput. Figure 4 scales the RIS aperture. As the surface grows, the legitimate SNR rises faster than the fixed-power jammer can follow, so antifragile benefits fade; the PS attack offers no gain beyond 128 elements. Across the source-aware scenario DRFM delivers the highest data rates, followed by AS and PS, consistent with mutual-information limits under fixed jammer power and position.
@@ V-B Source-Aware Eavesdropping
Figure 3: Throughput vs. JSR for optimal selected coding rate.
Baseline SNR = 10 dB
@@
V-C RIS-Aware-Eavesdropping
In RIS-aware eavesdropping scenario, we keep the RIS location, code rate, and modulation order identical to source-aware case. Figure 5 presents the reactive-jammer results under the same baseline SNR and adaptive coding settings used in Fig. 4. Here, every jammer achieves a noticeably higher throughput. This outcome is anticipated because the adversary receives the RIS-reflected waveform and thereby benefits from its beam-forming gain; consequently, the effective jamming SNR scales more rapidly as the JSR grows.
@@ VI Conclusion
This paper has presented the first study of antifragility in RIS-assisted wireless links under jamming attacks. We considered three attack models, i.e., DRFM, phase shifting (PS), amplitude shifting (AS). We proposed a receiver-side identification framework that was able to exploit spatial or temporal orthogonality to separate the desired and malicious waveforms, classifies the jammer type, and send the estimate back to the transmitter. A joint waveform-coding adaptation was devised: PSK/ASK remapping for AS and PS attacks, unchanged PSK for DRFM, and rate-adaptive Reed-Solomon coding that balances higher modulation orders against stronger error control. The results confirm that reconfigurable intelligent surfaces, when coupled with an antifragile design philosophy, can convert hostile interference into a throughput resource. This insight opens up a new design dimension for resilient wireless networks, where deliberate or accidental interference may be exploited rather than merely avoided.

### v2

@@ IV-D Signal, modulation and code adaptation
where $h_{L}$ denotes the cascaded legitimate channel form the destination to the RIS and $h_{E}$ and $h_{J}$ denote the eavesdropping and jamming channel based on Fig.1 $\begin{aligned} \gamma_{\mathrm{E}}&=\tfrac{P_{t}|h_{E}|^{2}}{\sigma_{E}^{2}},\quad\gamma_{\mathrm{J}}&=\tfrac{P_{j}|h_{J}|^{2}}{\sigma_{J}^{2}},\quad\gamma_{\mathrm{L}}&=\tfrac{P_{t}|h_{L}|^{2}}{\sigma_{L}^{2}}.\end{aligned}$ Since $(n,k)$ Reed-Solomon code can correct up to t=(nk)/2t=(n-k)/2 symbol errors, choosing the optimal code depends on the post-demodulation error profile. Thus, the measured SNR is used to compute the bit error rate (BER) for each adaptive modulation scheme, which is then converted into the residual symbol error rate $P^{\mathrm{RS}}_{\mathrm{res}}$ following the formulation in [Phung.2022].
@@
IV-E Antifragile gain
We denote the throughput in the baseline scenario by $T_{L}$ and under jamming attack by $T_{J}$ respectively. The optimal Reed-Solomon code is selected according to Eq. (15), which ensures that all symbol errors within each block are corrected. The throughput is then defined based on the code rate $R_{C}$, bandwidth $B$, modulation order $M$ and SNR index $i$ (Eq. 14):
@@ IV-E Antifragile gain
The system demonstrates antifragile gain only if its throughput of the paths with jammers surpasses its baseline, jammer-free throughput. To evaluate this gain across different SNRs, the jamming-to-signal ratio (JSR) is defined. It represents the ratio of jamming power $P_{J}$ to legitimate signal power $P_{L}$ between source and destination:
@@ V Numerical Evaluation
For simulations, we assume that RIS is positioned with respect to the receiver and transmitter at $d_{\mathrm{SR}}=18\penalty 10000\ \mathrm{m}$ and $d_{\mathrm{RD}}=7\penalty 10000\ \mathrm{m}$, respectively. We set the path-loss exponent to $\delta=2.7$ and the rate $\lambda_{R}=0.05$. The transmitter sends the legitimate signal with a fixed power of $20\penalty 10000\ \mathrm{dBm}$ on a $28\penalty 10000\ \text{GHz}$ carrier. Because the receiver’s signal-to-noise ratio (SNR) can fluctuate significantly depending on the size of the RIS, the jamming-to-signal ratio (JSR) is defined based on Eq. 17. This choice confines the jammer output to the interval $0\text{-}40\penalty 10000\ \text{dBm}$; an SNR-based definition would require substantially higher jammer powers to span the same JSR range at larger baseline SNRs. Adaptive modulation and coding (AMC) employs Reed-Solomon codes whose code rates vary from 0.70 (RS(178,255)(178,255)) to 0.94 (RS(240,255)(240,255)) based on Eq. 15. The modulation format is selected according to the attack detected, with orders of up to $M=64$. Results are reported separately for source-aware and RIS-aware eavesdropping paths, and antifragile gains are emphasized with light blue shading. In all simulations, the channel coefficients were iterated $200$ times, to obtain average values.
Figure 2 compares the three attacks with a fixed code rate. Without spatial orthogonality the jammer and desired signal overlap for half the frame; the transmitter shortens its burst, reducing payload and throughput (Eq. 16). Thus, antifragile gains appear only after raising the modulation order, i.e., at JSR = 3 dB for DRFM and 15 dB for AS. With full orthogonality and a higher baseline SNR, throughput exceeds the jammer, free reference from about -5 dB, DRFM allowing for the largest boost due to its high mutual information with the desired waveform. Figure 3 scales the RIS size. As the surface grows, the legitimate SNR increases faster than the fixed-power jammer can follow, so antifragile benefits fade; the PS attack offers no gain beyond 128 elements. Across the source-aware scenario DRFM delivers the highest data rates, followed by AS and PS, consistent with mutual-information limits under fixed jammer power and position.
@@
For RIS-aware eavesdropping path, we keep the RIS location, code rate, and modulation order identical to source-aware path. Figure 4 presents the reactive-jammer results under the same baseline SNR and adaptive coding settings used in Fig. 3. Here, every jammer achieves a noticeably higher throughput. This outcome is as expected because the adversary receives the RIS-reflected waveform and thereby benefits from its beam-forming gain. Consequently, the effective jamming SNR scales more rapidly as JSR grows.
@@ VI Conclusion
This paper presented for the first time a cross-layer network integration approach of RIS-assisted link antifragility under jamming attacks. The results confirm that RIS, when coupled with an antifragile design philosophy, can convert hostile interference into a throughput resource. The proposed antifragile behavior is the basis for cross-layer integration with network control mechanisms such as adaptive routing and link scheduling. This insight opens up a new design dimension for resilient wireless networks, where deliberate or accidental interference may be exploited rather than merely avoided.
Acknowledgment
This work was partially supported by the project MARE (Grant Agreement No 101191436), under Smart Networks and Services Joint Undertaking (SNS JU) EU Horizon program.